If
n(A) = 109, n(A ∪ B) = 140, and n(A ∩ B) = 44, find n(B).

Answer by ikleyn(52798)

.
If
n(A) = 109, n(A ∪ B) = 140, and n(A ∩ B) = 44, find n(B).
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

There is a remarkable formula

n(A U B) = n(A) + n(B) - n(A ∩ B).


Regarding this formula, see the lesson 

    - Counting elements in sub-sets of a given finite set

in this site.


By applying the formula, you have 

140 = 109 + n(B) - 44,

and you get

n(B) = 140 - 109 + 44 = 75.


Answer.  n(B) = 75.

Solved.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic "Miscellaneous word problems".